# Retrieval-Augmented Decision Transformer: External Memory for In-context RL

### v2

@@ 6 Conclusion
Existing in-context RL methods keep entire episodes in their context window, which is challenging as RL environments are typically characterized by long episodes and sparse rewards. To address this challenge, we introduce RA-DT, which employs an external memory mechanism to store past experiences and to retrieve experiences relevant for the current situation. RA-DT outperforms baselines on grid-worlds, while using only a fraction of their context length. While RA-DT improves average performance on holdout tasks on complex environments, it struggles to exhibit ICL, along with other in-context RL methods. Consequently, we illuminate the current limitations of in-context RL methods and discuss future directions. Finally, we release our datasets for Dark-Room, Dark Key-Door, MazeRunner, and Procgen, to facilitate future research on in-context RL.
Future Work. Besides the general directions discussed in Section 5, we highlight a number of concrete approaches to extend RA-DT. While we focus on ICL without relying on expert demonstrations, pre-filling the external memory with demonstrations may enable RA-DT to perform more complex tasks. This may be effective for robotics applications, where expert demonstrations are easy to obtain. Furthermore, end-to-end training of the retrieval component in RA-DT, similar to [Izacard et al., 2022], may result in more precise context retrieval and enhanced down-stream performance. Finally, we envision that modern recurrent architectures [Bulatov et al., 2022; Gu & Dao, 2023; Beck et al., 2024] as policy backbones may benefit RA-DT by maintaining hidden states across many episodes.
@@ Acknowledgments
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, and Leonardo at CINECA, Italy. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects Medical Cognitive Computing Center (MC3), INCONTROL-RL (FFG-881064), PRIMAL (FFG-873979), S3AI (FFG-872172), DL for GranularFlow (FFG-871302), EPILEPSIA (FFG-892171), AIRI FG 9-N (FWF-36284, FWF-36235), AI4GreenHeatingGrids (FFG- 899943), INTEGRATE (FFG-892418), ELISE (H2020-ICT-2019-3 ID: 951847), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01). We thank NXAI GmbH, Audi.JKU Deep Learning Center, TGW LOGISTICS GROUP GMBH, Silicon Austria Labs (SAL), FILL Gesellschaft mbH, Anyline GmbH, Google, ZF Friedrichshafen AG, Robert Bosch GmbH, UCB Biopharma SRL, Merck Healthcare KGaA, Verbund AG, GLS (Univ. Waterloo), Software Competence Center Hagenberg GmbH, Borealis AG, TÜV Austria, Frauscher Sensonic, TRUMPF and the NVIDIA Corporation.
@@ Appendix A Ethics Statement & Reproducibility
In recent years, there has been a trend in RL towards large-scale multi-task models that leverage offline pre-training. In this work, we broadly aim at building agents that can learn new tasks via ICL without the need for re-training or fine-tuning. Our goal is to reduce the need to provide entire past episodes in the agent’s context, by augmenting the agent with an external memory in combination with a retrieval component, similar to RAG in LLMs. We believe that multi-task agents of the near future will be able to perform a broad range of tasks, and that these agents will greatly benefit from RAG as used in RA-DT. The external memory component can enable agents to leverage information from in its own distant past or experiences from other agents. Such agents could have an immense impact on the global economy (e.g., as a source of inexpensive labour). As such, they do not come without risks and the potential for misuse. While we believe that our work can significantly impact the positive use of future agents, it is essential to ensure responsible deployment of future technologies.
We open-source the code-base used for our experiments, and release the datasets we generated. Both are available at: https://github.com/ml-jku/RA-DT. In addition, we provide further information on the environments/datasets, implementation including hyperparameter tables, and on our experiments in Appendices B, C, D, respectively.
@@ B.1 Dark-Room and Dark Key-Door
The Dark-Room environment is modelled after Morris-Watermaze, a classic experiment in behavioural neuroscience for studying spatial memory and learning in animals [D’Hooge & De Deyn, 2001]. We design our Dark-Room and Dark Key-Door environments in Minihack [Samvelyan et al., 2021], which is based on the NetHack Learning Environment [Küttler et al., 2020]. We construct grids of dimensions $10\times 10$, $20\times 20$ and $40\times 20$, as depicted in Figure 9. With increasing grid sizes, the task of locating the goal becomes harder as the number of possible positions in the grid grows (100, 400, 800). Therefore, we set the number of interaction steps per environment equal to the number of grid cells. Consequently, larger grids results in longer episodes and thus context lengths (e.g., 2400 for AD). The agent observes its own x-y position on the grid and can perform one of 5 actions at every interaction step (up, down, left, right, stay). Episodes start in the top left corner (0,0) and the agent is reset to the start position after every episode.
@@
In Dark-Room, the agent has to navigate to a randomly placed and invisible goal position. Therefore, the task space in Dark-Room environments is equal to the number of grid-cells (i.e., $100$ for $10\times 10$). The agent receives a reward for +1 for every step in the episode it is located in the goal position and 0 otherwise. As there are as many grid-cells as episode steps, the optimal strategy for solving the Dark-Room task is to use the first episode to visit every cell to find the hidden goal location. Once found, this knowledge can be exploited in upcoming trials.
@@
In contrast, in Dark Key-Door, there are two objects: a key and a goal state. Similar to Dark-Room, the key and goal position are randomly placed on the grid. The agent has to first pick up the invisible key and then find the invisible goal. Due to the presence of the two key events (picking up the key, finding the goal), the task space is combinatorial in the number of grid-cells (i.e., $10000100^{2}=$ 10000 for $10\times 10$). This makes the Dark Key-Door more challenging than the Dark-Room task, especially as the grid-size becomes larger.
(a) Dark-Room 10$\times$10
(b) Dark-Room 20$\times$20
(c) Dark-Room 40$\times$20
Figure 7: Average performances of the source algorithm, PPO, on 80 train tasks for Dark-Room (a) 10$\times$10, (b) 20$\times$20, and (c) 40$\times$20. For (a), (b) we train PPO on individual tasks for 100K environment steps. For (c), we train for 200K environment steps to take the longer episode lengths into account. We evaluate the agents after every 10K steps. Curves show the mean reward achieved (+ 95% CI) across the 80 train tasks.
@@
Training Dataset. For both Dark-Room and Dark Key-Door, we generate training datasets for 80 randomly assigned goals or key-goal combinations. We use PPO [Schulman et al., 2017] to generate 100K environment transitions per goal location for $10\times 10$ and $20\times 20$ grids and 200K environment transitions for the largest grid. Therefore, the total number of transitions across datasets is 8M for $10\times 10$ and $20\times 20$ grids and 16M for $40\times 20$.
@@
We train PPO with standard hyperparameter settings in stable-baselines3 [Raffin et al., 2021] using a learning rate of $3e^{-4}$, batch size of $64$, number of steps between updates of $2048$, number of update epochs $10$ and entropy coefficient of $0.01$. For $20\times 20$ and $40\times 20$ grids, we increase the number of update epochs to $30$ and the entropy coefficient of to $0.1$ for $40\times 20$. We store all generated transitions of PPO for our datasets. Consequently, the final datasets contain a mixture of suboptimal or exploratory, and optimal or exploitative behaviour.
@@
Source Algorithm Performance. We show average learning curves across all task-specific PPO agents on the 80 training tasks for all grid-sizes in Figures 7 and 8 for Dark-Room and Dark Key-Door, respectively. For the $10\times 10$ grids, the average performance converges towards optimal performance. However, on the larger grid sizes, the performances are below the optimum. This is because it takes the agent longer to discover and collect successful episodes by initially random environment interaction as the grids become larger.
(a) Dark Key-Door 10$\times$10
(b) Dark Key-Door 20$\times$20
(c) Dark Key-Door 40$\times$20
Figure 8: Average performances of the source algorithm, PPO, on 80 train tasks for Dark Key-Door (a) 10$\times$10, (b) 20$\times$20, and (c) 40$\times$20. For (a), (b) we train PPO on individual tasks for 100K environment steps. For (c), we train for 200K environment steps. We evaluate the agents after every 10K steps. Curves show the mean reward achieved (+ 95% CI) across the 80 train tasks.
(a) Room 10$\times$10
(b) Key-Door 10$\times$10
(c) Room 20$\times$20
(d) Room 40$\times$20
Figure 9: Mini-grid environments. In Dark-Room, the agent is located in a room and has to navigate to an invisible goal location. We use grid-sizes (a) 10$\times$10, (b) 20$\times$20 and (c) 40$\times$20 for our experiments. In (b) Dark-KeyDoor, the agent has to pick up an invisible key, then navigate to the invisible goal location. Agents only observe their current x-y coordinate on the grid. Reward of +1 is obtained in every step the agent is situated in the goal state, +1 for picking up the key.
@@ B.2 MazeRunner
MazeRunner was introduced by [Grigsby et al., 2023] and inspired by the Memory Maze environment [Pasukonis et al., 2022]. The agent is located in a 15$\times$15 procedurally-generated maze and has to navigate to a sequence of one, two, or three goal locations in the right order (see Figure 10). Similar to Dark-Room environments, MazeRunner is partially observable and exhibits sparse rewards. The agent observes a Lidar-like 6-dimensional representation of the state that contains 4 continuous values that measure the distance from the agent’s location to the nearest wall, and the x-y coordinates of the agent’s position in the grid. The action-space is 4-dimensional (up, down, left, right). A reward of +1 is obtained when reaching the currently active goal state in the goal sequence. Therefore, the total achievable reward is equal to the number of goal states. Episodes last for a maximum of 400 steps or terminate early, if all goal locations have been reached. After every episode, the agent (gray box in Figure 10) is reset to the origin location. During evaluation, we allow for 30 ICL trials, which amounts to 12K environment steps in total.
@@
Figure 10: Maze-Runner environments introduced by Grigsby et al. [2023]. In Maze-Runner, the agent is located in a procedurally generated $15\times 15$ maze and has to navigate to (a) one, (b) two or (c) goal locations in pre-specified order. The agent receives a reward of +1 for reaching a goal. Episodes last for a maximum of 400 steps, or terminate early if all goal locations have been visited.
Training Dataset. The procedural-generation of the maze and selection of the number of goals is controlled by setting the environment seed. We use PPO to generate 100K environment interactions for 100 procedurally-generated mazes, and record the entire replay buffer, which amounts to 10M transitions in total. We found it necessary, to equip the task-specific PPO agents with an LSTM [Hochreiter & Schmidhuber, 1997] policy. Without the LSTM, agents hardly make progress for some mazes, especially if the maze contains two or three goal locations. For this reason, we first generate data for more than 100 mazes and select the first 100 seeds, where the average reward at the end of training is $>0.25$. This results in a set of seeds in $[0,120]$ Otherwise, we use standard hyperparameter settings as provided in stable-baselines3.
Source Algorithm performance. We show the average learning curves over all 100 task-specific PPO agents in Figure 11. On average, the agents receive a reward of $\approx 1$ over all mazes. This average include environments with one, two or three goals. We provide further dataset statistics for MazeRunner with the corresponding dataset release.
Figure 11: Learning curves for data-collection runs on all 100 mazes on Maze-Runner 15$\times$15 environments with PPO-LSTM as source algorithm. We train for 100K environment steps on each maze and report the mean reward achieved (+ 95% CI).
@@ B.3 Meta-World
The Meta-World benchmark [Yu et al., 2020a] consists of 50 challenging robotics tasks, such as opening/closing a window, using a hammer, or pressing buttons. All tasks in Meta-World use a Sawyer robotic arm simulated using the MuJoCo physics engine [Todorov et al., 2012]. The observations and actions are 39-dimensional and 6-dimensional continuous vectors, respectively. As all tasks share the robotic arm, the state, and action spaces remain constant across tasks. All actions are in range $[-1,1]$. The reward functions are dense and based on distances to the goal locations (exact reward-definitions are provided in Yu et al. [2020a]). Similar to Wolczyk et al. [2021] and Schmied et al. [2024], we limit the episode lengths to 200 interactions. We follow Yu et al. [2020a] and split the 50 Meta-World tasks into 45 training tasks (ML45) and 5 evaluation tasks (ML5). During evaluation, we use deterministic environment resets after episodes, i.e., objects and goal positions are reset to their original state. Furthermore, we mask-out the goal positions in the state vector, which forces agents to adapt during environment interaction. Agents are given 30 ICL trials during evaluation. The 5 evaluation tasks are:
@@
Training Dataset. For our Meta-World experiments, we leverage the datasets released by Schmied et al. [2024]. The datasets contain 2M transitions per task, which amounts to 90M transitions across all ML45 training tasks. The data was generated with randomized object and goal positions after every episode.
@@ B.4 DMControl
DMControl contains 30 different robotic tasks with different robot morphologies [Tassa et al., 2018]. Similar to prior work [Hafner et al., 2019; Schmied et al., 2024], we select 16 of these 30 tasks and split them into 11 training (DMC11) and 5 evaluation tasks (DMC5). The DMC11 training tasks are:
@@
States and actions in DMControl are continuous vectors. As DMControl contains different robot morphologies, the state, and action spaces vary considerably across tasks ($3\leq|\mathcal{S}|\leq 24$, $1\leq|\mathcal{A}|\leq 6$). All actions in DMControl are bounded by $[-1,1]$. Episodes last for 1000 environment steps and per time-step a maximum reward of +1 can be achieved, which results in a maximum reward of 1000 per episode. Agents are given 30 ICL trials per task during evaluation, which results in 30K steps for a single evaluation run.
Training Dataset. As for Meta-World, we leverage the datasets released by Schmied et al. [2024]. The datasets contain 1M transitions per task, which amounts to 11M transitions used for training across all DMC11 tasks. We refer to Schmied et al. [2024] for further dataset statistics on DMControl and Meta-World.

### v3

@@ 6 Conclusion
Existing in-context RL methods keep entire episodes in their context window, which is challenging as RL environments are typically characterized by long episodes and sparse rewards. To address this challenge, we introduce RA-DT, which employs an external memory mechanism to store past experiences and to retrieve experiences relevant to the current situation. The external memory component in RA-DT enables our agents to leverage information from their own distant past or even experiences from other agents. The ability to access experiences from the distant past can be particularly relevant for scenarios in which agents learn over extended time horizons, such as continual learning setups. RA-DT outperforms baselines on grid-worlds while using only a fraction of their context length. While RA-DT improves average performance on holdout tasks in complex environments, our approach struggles to exhibit ICL, along with other in-context RL methods. Consequently, we illuminate the current limitations of in-context RL methods and discuss future directions. Finally, we release our datasets for Dark-Room, Key-Door, MazeRunner, and Procgen, to facilitate future research in-context RL research.
Future Work. Besides the general directions discussed in Section 5, we highlight several concrete approaches to extend RA-DT. While we focus on ICL without relying on expert demonstrations, pre-filling the external memory with demonstrations may enable RA-DT to perform more complex tasks. This approach may be effective for robotics applications, where expert demonstrations are easy to obtain. In contrast, our current approach relies solely on the self-improvement abilities of the trained agent. Furthermore, end-to-end training of the retrieval component in RA-DT, similar to (Izacard et al., 2022), may lead to more precise context retrieval and improved downstream performance. Finally, we envision that modern recurrent architectures (Bulatov et al., 2022; Gu & Dao, 2023; Beck et al., 2024) as policy backbones may strongly benefit RA-DT by maintaining hidden states across many episodes.
@@ Acknowledgments
We acknowledge EuroHPC Joint Undertaking for awarding us access to Karolina at IT4Innovations, Czech Republic, MeluXina at LuxProvide, Luxembourg, and Leonardo at CINECA, Italy. The ELLIS Unit Linz, the LIT AI Lab, the Institute for Machine Learning, are supported by the Federal State Upper Austria. We thank the projects FWF AIRI FG 9-N (10.55776/FG9), AI4GreenHeatingGrids (FFG- 899943), Stars4Waters (HORIZON-CL6-2021-CLIMATE-01-01), FWF Bilateral Artificial Intelligence (10.55776/COE12). We thank NXAI GmbH, Audi AG, Silicon Austria Labs (SAL), Merck Healthcare KGaA, GLS (Univ. Waterloo), TÜV Holding GmbH, Software Competence Center Hagenberg GmbH, dSPACE GmbH, TRUMPF SE + Co. KG.
@@ Appendix A Ethics Statement & Reproducibility
In recent years, there has been a trend in RL towards large-scale multi-task models that leverage offline pre-training. In this work, we broadly aim at building agents that can learn new tasks via ICL without the need for re-training or fine-tuning. Our goal is to reduce the need to provide entire past episodes in the agent’s context by augmenting the agent with an external memory in combination with a retrieval component, similar to RAG in LLMs. We believe that multi-task agents of the near future will be able to perform a broad range of tasks, and that these agents will greatly benefit from RAG as used in RA-DT. The external memory component can enable agents to leverage information from their own distant past or experiences from other agents. Such agents could have an immense impact on the future of work (e.g., as a source of inexpensive labour). As such, they do not come without risks and the potential for misuse. While we believe that our work can significantly impact the positive use of future agents, it is essential to ensure the responsible deployment of future technologies.
We open-source the codebase used for our experiments and release the datasets we generated  [^2] GitHub: https://github.com/ml-jku/RA-DT. In addition, we provide further information on the environments/datasets, implementation including hyperparameter tables, and on our experiments in Appendices B, C, D, respectively.
@@ B.1 Dark-Room and Dark Key-Door
The Dark-Room environment is modelled after Morris-Watermaze, a classic experiment in behavioural neuroscience for studying spatial memory and learning in animals (D’Hooge & De Deyn, 2001). We design our Dark-Room and Dark Key-Door environments in Minihack (Samvelyan et al., 2021), which is based on the NetHack Learning Environment (Küttler et al., 2020). We construct grids of dimensions $10\times 10$, $20\times 20$, and $40\times 20$, as depicted in Figure 9. With increasing grid sizes, the task of locating the goal becomes harder as the number of possible positions in the grid grows (100, 400, 800). Therefore, we set the number of interaction steps per environment equal to the number of grid cells. Consequently, larger grids result in longer episodes and thus context lengths (e.g., 2400 for AD). The agent observes its own x-y position on the grid and can perform one of 5 actions at every interaction step (up, down, left, right, stay). Episodes start in the top left corner (0,0), and the agent is reset to the start position after every episode.
@@
In Dark-Room, the agent has to navigate to a randomly placed and invisible goal position. Therefore, the task space in Dark-Room environments is equal to the number of grid-cells (i.e., $100$ for $10\times 10$). The agent receives a reward of +1 for every step in the episode if it is located in the goal position and 0 otherwise. As there are as many grid-cells as episode steps, the optimal strategy for solving the Dark-Room task is to use the first episode to visit every cell to find the hidden goal location. Once found, this knowledge can be exploited in upcoming trials.
@@
In contrast, in Dark Key-Door, there are two objects: a key and a goal state. Similar to Dark-Room, the key and goal positions are randomly placed on the grid. The agent has to first pick up the invisible key and then find the invisible goal. Due to the presence of the two key events (picking up the key, finding the goal), the task space is combinatorial in the number of grid-cells (i.e., $100^{2}=10000$ for $10\times 10$). This makes the Dark Key-Door more challenging than the Dark-Room task, especially as the grid size becomes larger.
(a) Dark-Room $\times10$
(b) Dark-Room $\times20$
[…]
@@
Training Dataset. For both Dark-Room and Dark Key-Door, we generate training datasets for 80 randomly assigned goals or key-goal combinations. We use PPO (Schulman et al., 2017) to generate 100K environment transitions per goal location for $10\times 10$ and $20\times 20$ grids and 200K environment transitions for the largest grid. Therefore, the total number of transitions across datasets is 8M for $10\times 10$ and $20\times 20$ grids and 16M for $40\times 20$.
@@
We train PPO with standard hyperparameter settings in stable-baselines3 (Raffin et al., 2021) using a learning rate of $3e^{-4}$, batch size of $64$, number of steps between updates of $2048$, number of update epochs $10$, and entropy coefficient of $0.01$. For $20\times 20$ and $40\times 20$ grids, we increase the number of update epochs to $30$ and the entropy coefficient to $0.1$ for $40\times 20$. We store all generated transitions of PPO for our datasets. Consequently, the final datasets contain a mixture of suboptimal or exploratory and optimal or exploitative behaviour.
@@
Source Algorithm Performance. We show average learning curves across all task-specific PPO agents on the 80 training tasks for all grid-sizes in Figures 7 and 8 for Dark-Room and Dark Key-Door, respectively. For the $10\times 10$ grids, the average performance converges towards optimal performance. However, on the larger grid sizes, the performance remains below the optimum. This is because it takes the agent longer to discover and collect successful episodes by initially random environment interaction as the grids become larger.
(a) Dark Key-Door $\times10$
(b) Dark Key-Door $\times20$
[…]
(a) Room $\times10$
(b) Key-Door $\times10$
(c) Room $\times20$
[…]
@@ B.2 MazeRunner
MazeRunner was introduced by (Grigsby et al., 2023) and inspired by the Memory Maze environment (Pasukonis et al., 2022). The agent is located in a $\times15$ procedurally-generated maze and has to navigate to a sequence of one, two, or three goal locations in the right order (see Figure 10). Similar to Dark-Room environments, MazeRunner is partially observable and exhibits sparse rewards. The agent observes a Lidar-like 6-dimensional representation of the state that contains 4 continuous values that measure the distance from the agent’s location to the nearest wall, and the x-y coordinates of the agent’s position in the grid. The action space is 4-dimensional (up, down, left, right). A reward of +1 is obtained when reaching the currently active goal state in the goal sequence. Therefore, the total achievable reward is equal to the number of goal states. Episodes last for a maximum of 400 steps or terminate early if all goal locations have been reached. After every episode, the agent (gray box in Figure 10) is reset to the origin location. During evaluation, we allow for 30 ICL trials, which amounts to 12K environment steps in total.
@@
Figure 10: Maze-Runner environments introduced by Grigsby et al. (2023). In Maze-Runner, the agent is located in a procedurally generated $15\times 15$ maze and has to navigate to (a) one, (b) two, or (c) goal locations in pre-specified order. The agent receives a reward of +1 for reaching a goal. Episodes last for a maximum of 400 steps, or terminate early if all goal locations have been visited.
Training Dataset. The procedural generation of the maze and selection of the number of goals is controlled by setting the environment seed. We use PPO to generate 100K environment interactions for 100 procedurally-generated mazes, and record the entire replay buffer, which amounts to 10M transitions in total. We found it necessary to equip the task-specific PPO agents with an LSTM (Hochreiter & Schmidhuber, 1997) policy. Without the LSTM, agents hardly make progress for some mazes, especially if the maze contains two or three goal locations. For this reason, we first generate data for more than 100 mazes and select the first 100 seeds, where the average reward at the end of training is >0.25>0.25. This results in a set of seeds in $[0,120]$ Otherwise, we use standard hyperparameter settings as provided in stable-baselines3.
Source Algorithm performance. We show the average learning curves over all 100 task-specific PPO agents in Figure 11. On average, the agents receive a reward of $\approx 1$ over all mazes. This average includes environments with one, two, or three goals. We provide further dataset statistics for MazeRunner with the corresponding dataset release.
Figure 11: Learning curves for data-collection runs on all 100 mazes on Maze-Runner $\times15$ environments with PPO-LSTM as source algorithm. We train for 100K environment steps on each maze and report the mean reward achieved (+ 95% CI).
@@ B.3 Meta-World
The Meta-World benchmark (Yu et al., 2020a) consists of 50 challenging robotics tasks, such as opening/closing a window, using a hammer, or pressing buttons. All tasks in Meta-World use a Sawyer robotic arm simulated using the MuJoCo physics engine (Todorov et al., 2012). The observations and actions are 39-dimensional and 6-dimensional continuous vectors, respectively. As all tasks share the robotic arm, the state and action spaces remain constant across tasks. All actions are in the range [1,1][-1,1]. The reward functions are dense and based on distances to the goal locations (exact reward definitions are provided in Yu et al. (2020a)). Similar to Wolczyk et al. (2021) and Schmied et al. (2024), we limit the episode lengths to 200 interactions. We follow Yu et al. (2020a) and split the 50 Meta-World tasks into 45 training tasks (ML45) and 5 evaluation tasks (ML5). During evaluation, we use deterministic environment resets after episodes, i.e., objects and goal positions are reset to their original state. Furthermore, we mask out the goal positions in the state vector, which forces agents to adapt during environment interaction. Agents are given 30 ICL trials during evaluation. The 5 evaluation tasks are:
@@
Training Dataset. For our Meta-World experiments, we leverage the datasets released by Schmied et al. (2024). The datasets contain 2M transitions per task, which amounts to 90M transitions across all ML45 training tasks. The data was generated with randomized object and goal positions after every episode.
@@ B.4 DMControl
DMControl contains 30 different robotic tasks with different robot morphologies (Tassa et al., 2018). Similar to prior work (Hafner et al., 2019; Schmied et al., 2024), we select 16 of these 30 tasks and split them into 11 training (DMC11) and 5 evaluation tasks (DMC5). The DMC11 training tasks are:
@@
States and actions in DMControl are continuous vectors. As DMControl contains different robot morphologies, the state and action spaces vary considerably across tasks ($3\leq|\mathcal{S}|\leq 24$, $1\leq|\mathcal{A}|\leq 6$). All actions in DMControl are bounded by [1,1][-1,1]. Episodes last for 1000 environment steps, and per time-step, a maximum reward of +1 can be achieved, which results in a maximum reward of 1000 per episode. Agents are given 30 ICL trials per task during evaluation, which results in 30K steps for a single evaluation run.
Training Dataset. As for Meta-World, we leverage the datasets released by Schmied et al. (2024). The datasets contain 1M transitions per task, which amounts to 11M transitions used for training across all DMC11 tasks. We refer to Schmied et al. (2024) for further dataset statistics on DMControl and Meta-World.